OpenStudy (anonymous):

Simplify (4.5)(5)(−2). 45 4.5 −4.5 −45

OpenStudy (anonymous):

@iGreen @rockinhood

OpenStudy (igreen):

Same question..?

OpenStudy (anonymous):

sorry no

OpenStudy (anonymous):

Simplify negative 4 over 5 ÷ 3 over negative 2. negative 8 over 15 negative 6 over 5 6 over 5 8 over 15

OpenStudy (anonymous):

:)

OpenStudy (igreen):

\(\sf\dfrac{4}{5} \div -\dfrac{3}{2} \rightarrow \dfrac{4}{5} \times -\dfrac{2}{3}\)

OpenStudy (igreen):

We flip the 2nd fraction and multiply when we divide.

OpenStudy (anonymous):

WHOA BOY HOLD UP

OpenStudy (anonymous):

-0.533333333333333333 for the first half

OpenStudy (anonymous):

for both halves acctually

OpenStudy (anonymous):

@rockinhood @iGreen ??

OpenStudy (igreen):

Just multiply the numerators and denominators together. \(\sf\dfrac{4}{5} \times -\dfrac{2}{3} \rightarrow \dfrac{4\times-2}{5\times3}\rightarrow ~ \! ?\)

OpenStudy (anonymous):

I SUCK AT THAT

OpenStudy (anonymous):

I CANT DO IT :(

OpenStudy (igreen):

You suck at multiplying? 4 * -2 = ? 5 * 3 = ?

OpenStudy (anonymous):

FRACTIONS

OpenStudy (igreen):

No, we're just multiplying whole numbers right now..when we multiply fractions together, we just multiply the tops together and the bottoms together.

OpenStudy (igreen):

So can you multiply 4 * -2 and 5 * 3?

OpenStudy (anonymous):

-8 and 15

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Yes, so our answer is \(\sf-\dfrac{8}{15}\).

OpenStudy (anonymous):

okie

OpenStudy (anonymous):

NUTHER Q?

OpenStudy (igreen):

Sure

OpenStudy (anonymous):

new one !!